Consider a one-year, 10-percent coupon bond with a face value of $1,000 issued by a private corporation.

The one-year risk-free rate is 10 percent.

The corporation has hit on hard times, and the consensus is that there is a 20 percent probability that it will default on its bonds.

If an investor were willing to pay at most $775 for the bond, is that investor risk-neutral or risk averse?
Answer #1

In case a bond is risk-free, then it would pay off $1,100 in one year's time that includes $1000 as principal and $100 as coupon.

If there is a 20% risk of default, then the expected value will be

Expected value = 1100*0.80+0*0.20 =$880

The present value of expected value in one year's time = 880/1.1 = $800 the price which the risk-neutral investor will be agreeing to pay.

If an investor were willing to pay at most $775 for the bond, it means he is willing to pay lower then the price agreed by the risk-neutral player, that implies that investor wants to get compensation for taking the risk associated with the bond, so he is risk-averse.